OpenStudy (emilyjones284):

Using u-sub 5csc^2(3t) dt

OpenStudy (danjs):

let u = 3t du = 3 dt

OpenStudy (danjs):

recall the derivative of tan(x) is csc^2(x)

OpenStudy (danjs):

\[\frac{ 5 }{ 3 } \int\limits \csc ^{2}(u) du\]

OpenStudy (danjs):

dt = 1/3 du

OpenStudy (danjs):

\[\frac{ -5 }{ 3 \tan(u) }\] sub back in for u = 3x

OpenStudy (danjs):

been a few years since calc 1. haha

OpenStudy (emilyjones284):

Thank you! I understand now (: